OpenStudy (anonymous):

The average surface pressure on Mars is 0.80 kPa and the acceleration of gravity is only 3.7 m/s2. With a barometer on Mars, what is the height of the column of mercury, assuming that mercury is liquid with a density of 13.6 × 103 kg/m3?

sam (.sam.):

Do you remember\[P=\rho g h\]?

OpenStudy (anonymous):

I'm not sure, about unit

sam (.sam.):

Pressure will be in pascals \[800=(13.6)(3.7)(h) \\ \\ h=\frac{800}{13.6 \times 3.7}\]

sam (.sam.):

\[800=(13.6)(3.7)(h) \\ \\ h=\frac{800}{(13.6\times 10^3) \times 3.7} \huge ^{*}\]

OpenStudy (anonymous):

Ok thank u so much

OpenStudy (anonymous):

We don't care about unit?

OpenStudy (anonymous):

The answers is 16mm

sam (.sam.):

Yeah

sam (.sam.):

The units cancels out giving meters, which is height.

sam (.sam.):

Is 16mm the actual answer or that's what you found out?

OpenStudy (anonymous):

Nope, that is my teacher answer

sam (.sam.):

ok we got that

sam (.sam.):

\[800Pa=(13.6\times 10^{3}kgm^{-3})(3.73.7ms^{-2})(h)\] Pascal has units \(\Large kgm^{-1}s^{-1}\) \[h=\frac{800kgm^{-1}s^{-1}}{(13.6\times 10^3kgm^{-3}) \times 3.7ms^{-2}}\] \[h=0.0159m=0.016m=16mm\]

OpenStudy (anonymous):

Thanks